4. Write two equations to depict the amphitropic nature of water.


(b) define the following: (I) pH and (II) pKa


(c).For a reaction between a weak acids and water as shown, prove that pH =pKa -log [acid/base].

HB(acid) + H2O = H3O+ + B-


(d) List the factors that affect the solubility of a substances.


(e) Calculate the solubility in grams per litre of silver sulphide if the Ksp of Ag2S is 6.3 x 10-50

1
Expert's answer
2022-02-25T05:28:03-0500

A) The pH is a measure of the concentration of hydrogen ions in an aqueous solution. ... Essentially, pKa tells you what the pH needs to be in order for a chemical species to donate or accept a proton. The relationship between pH and pKa is described by the Henderson-Hasselbalch equation

B) The procedure for calculating the pH of a solution of a weak base is similar to that of the weak acid in the example. However, the variable x will represent the concentration of the hydroxide ion. The pH is found by taking the negative logarithm to get the pOH, followed by subtracting from 14 to get the pHpH = pKa + log[A+]/[HA]


Modified... for weak bases

pOH = pKb + log[HB]/[B-]

pOH = pKb + log[Acid]/[Base]


Let's test it out dude....

If the pH = 3.8, then the pOH = 14 - 3.8

pOH = 10.2

Also, since the percent ionization of HB was 67%, that means there is 67% B- and 33% HB in solution.

Use these values in the equation to solve for pKb

pOH = pKb + log[HB]/[B-]

10.2 = pKb + log[0.33]/[0.67]

10.2 = pKb - 0.31

pKb = 10.2 + 0.31

pKb = 10.51


*Fingers crossed Blake...

One last thing.

HB is the acid,

H2O is the base

B- is the conjugate base

H3O+ is the conjugate acid